OpenStudy (anonymous):

Factor completely 81x2 - 4 Choose one answer. a. (2x - 9)(2x - 9) b. (2x - 9)(2x + 9) c. (9x - 2)(9x - 2) d. (9x - 2)(9x + 2)

OpenStudy (anonymous):

I got D

OpenStudy (anonymous):

sound right?

OpenStudy (anonymous):

a² - b² = (a + b)(a - b) so in this case 81x² - 4 = (9x + 2) (9x - 2)

OpenStudy (ryan123345):

Yes, D is correct! :)